OpenStudy (anonymous):

Neal opened a savings account 11 years ago with a deposit of $2,335.69. The account has an interest rate of 4.3% compounded daily. How much interest has Nelson earned? I ended up getting an extremely high number i'm not really sure what i did wrong

OpenStudy (anonymous):

i put the formula as A=2,335.69(1+.043/11)^4015

OpenStudy (ranga):

It should be A = 2335.69(1 + .043/365)^(365 * 11)

OpenStudy (anonymous):

Watch your n inside the parenthesis. You have n = 11 inside the parenthesis and n = 365 in the exponent. Figure out which one it should be, then make sure all the "n"s are the same.

OpenStudy (anonymous):

so then wouldnt the (365*11) turn into 4015?

OpenStudy (ranga):

Your exponent is 4015. But your mistake is in .043/11 but it should be .043/365

OpenStudy (anonymous):

ohhhhhhh! okay let me try again

OpenStudy (anonymous):

yeah i did it wrong i got 1

OpenStudy (ranga):

wrong now?

OpenStudy (anonymous):

I did uhm .043/365 + 1 and got 1

OpenStudy (ranga):

You have to go several decimal places in these kinds of calculations. Don't round off until the end.

OpenStudy (anonymous):

okay well i got 1.00011781 then

OpenStudy (anonymous):

so after exponent got 1.60476814741

OpenStudy (anonymous):

ooo okay i got the answer! thank you!

OpenStudy (anonymous):

What was the answer you got?

OpenStudy (anonymous):

$3,748.21

OpenStudy (ranga):

Cool. Did you subtract the principal from the maturity amount to find the interest?

OpenStudy (ranga):

Interest earned = A - P = $3,748.21 - $2,335.69 = ?

OpenStudy (anonymous):

That was the purpose of my question. That number you got at the end was the final amount in the account. To find the interest earned, you have to subtract the original principal from the new balance.

OpenStudy (anonymous):

oh crap i didn't i thought that you normally have to add that amount to the biggening amount to get the ending?

OpenStudy (anonymous):

No. The formula you just used doesn't give you the interest, it gives you the new amount of money in the account. So if you started with 2335.69, after 11 years, you'll have 3748.21. The interest earned is the amount of money that was added to the principal to get the new amount, or the difference between the two amounts. That formula doesn't give you the interest. It gives you the new amount.

OpenStudy (ranga):

For simple interest the formula is P * r * t. You calculate the interest and then you add it to the principal to get the ending balance. The compound interest formula we used gives the ending balance. So you will have to subtract the beginning amount to figure the interest paid to the account.

OpenStudy (anonymous):

the answer is not 3748.21